how do I find the sum of consecutive integers from 17 to 97? what is the formula for it?
how to find the sum of consecutive odd integers from 17 to 97?

Answer by richwmiller(17219)

an = a + (n - 1)d
solve for n
97 = 17 + (n - 1)1
n=81
97 = 17 + (n - 1)21
n=41 odd
plug in n below
S = ½(a + an)n
S = ½(17 + 97)81=4617
S = ½(17 + 97)41=2331